# Plain Transformers Can be Powerful Graph Learners

### v2

@@ 3.4 Sinusoidal PE Enhancement (SPE)
As discussed in Zhang et al., 2023b ; Zhang et al., 2024b , the finer-grained the information provided through graph positional encodings (PEs), the stronger distinguishability of graph structure GD-GNNs can achieve. However, due to the spectral bias of neural networks Rahaman et al., (2019), MLPs prioritize learning the low-frequency modes and thus lose detailed information stored in PE.
To mitigate this issue, motivated by NeRF Mildenhall et al., (2020), we add an extra sinusoidal encoding on top of the RRWP and process it with a simple MLP. The sinusoidal encoding is applied to each channel of $\mathbf{p}^{\prime}_{ij}$ in an elementwise fashion:
@@
where $\mathtt{SPE}:\mathbb{R}\to\mathbb{R}^{1+2S}$ with $S\in\mathbb{Z}^{+}$ different bases. For notational conciseness, we use $\mathtt{SPE}(\mathbf{p}^{\prime}_{ij})\in\mathbb{R}^{K+2SK}$ to denote the application of $\mathtt{SPE}$ to all channels of $\mathbf{p}^{\prime}_{ij}$ and concatenate the outputs. The PE stem (Eq. 8) becomes
@@ 3.5 Theoretical Expressivity of PPGT
The theoretical expressivity of PPGT can be analyzed with the GD-WL framework Zhang et al., 2023b . Since it is not proposed in our work, we provide further discussion of it in Appx. E.1.
@@ 4 Relationship with Previous Work
Graph Transformers Motivated by the success of Transformers in other domains, researchers have strived to migrate Transformers to graph learning to address the limitations of MPNNs. Although the naive migration of Transformer to graph learning did not work well Dwivedi and Bresson, (2021), several recent works have achieved considerable success when constructing graph Transformers, from theoretical expressivity analysis Zhang et al., 2023b ; Zhang et al., 2024b to outstanding empirical performance Ying et al., (2021); Rampášek et al., (2022); Ma et al., (2023). However, during the development of these graph Transformers, there has been a gradual but substantial deviation from the plain Transformers widely used in other domains. For example, SAN Kreuzer et al., (2021) introduces dual-attention mechanisms with local and global aggregations; K-Subgraph SAT Chen et al., (2022) introduce MPNNs into attention mechanisms to compute attention scores; GraphGPS Rampášek et al., (2022) heavily relies on the MPNNs within its hybrid Transformer architecture; EGT Hussain et al., (2022) introduces gated mechanisms and edge-updates inside the attention; GRIT Ma et al., (2023) incorporates a complicated conditional MLP-based attention mechanism (shown in Appendix. D.1). These deviations prevent the easy adoption of Transformer training advances and obscure the potential unification of cross-modality foundation models. Among these graph Transformers, the graphormer-series Ying et al., (2021); Luo et al., (2022); Zhang et al., 2023b retain architectures that are closest to plain Transformers; unfortunately, the result is a substantial gap in empirical performance and empirical expressivity to the best-performing graph Transformers Rampášek et al., (2022); Ma et al., (2023). In this work, we analyze the limitations of plain Transformers for learning on graphs and propose simple but effective designs to allow plain transformers to reach outstanding empirical performance as well as remarkable expressivity.
@@
Other Attention Mechanisms Besides SDP attention, there are other attention variants in use. Bahdanau et al., (2015) introduced the earliest content-based attention mechanism for recurrent neural networks (RNNs) based on an MLP, which is more computational and memory costly. Swin-Transformer-V2 Liu et al., (2022) uses cosine-similarity to compute attention scores for better stability, but neglects the magnitude information. Kim et al., (2021) propose the use of the negative of the square $L_{2}$-distance, with tied query-key projection weights, for maintaining Lipchitz continuity of Transformers. Our attention mechanism, although based on $L_{2}$ attention, is further simplified and adjusted in order to maintain alignment with SDP attention.
@@
Additional related work We discuss additional related work, including MPNNs, graph positional/structural encoding, higher-order GNNs, and subgraph GNNs in Appx. D.
@@ 5.1 Empirical Expressivity on Graph Isomorphism
To better understand the expressivity of PPGT, we evaluate our model on the BREC benchmark Wang and Zhang, (2024), a comprehensive dataset for measuring the empirical expressive power of GNNs w.r.t. graph isomorphism, with graph-pairs from 1-WL to 4-WL-indistinguishable.
@@
[3]. Mismatch between expressivity and real-world benchmark performance: The stronger theoretical/empirical expressivity is not completely reflected in real-world benchmark performance. For example, subgraph GNNs and/or K-WL GNNs with stronger expressivity (e.g., SSWL+, I2GNN, N2GNN) reach inferior performance compared to GRIT and PPGT on the ZINC benchmark.
[4]. Going beyond GD-WL?: As previously discussed, the design of graph PE, a.k.a., pseudo-coordiantes, can be extended beyond the distance/affinity of graphs. With this in mind, we conduct a exploratory demo called I2GNN+PPGT, which uses I2GNN to generate additional positional encodings for PPGT. The empirical expressivity is further improved to 76%, outperforming the standalone I2GNN and PPGT, surpassing 3-WL and reaching the top performance among the methods compared. This demonstration hints that plain Graph Transformers can potentially surpass GD-WL and achieve greater expressive power purely through enhanced positional encoding designs.
@@ 5.2 Benchmarking PPGT on Real-world Benchmarks
The specifications, details, and references for the baseline methods are provided in Appx. A.5.
@@
We conduct a general evaluation of our proposed PPGT on five datasets from Benchmarking GNNs (Dwivedi et al., 2022a, ): ZINC, MNIST, CIFAR10, PATTERN, and CLUSTER, and summarize the results in Table 2. We observe that our model obtains the best mean performance for all five datasets, outperforming various MPNNs, non-MPNN GNNs, and existing graph Transformers. These results showcase the effectiveness of PPGT for general graph learning with a plain Transformer architecture.
Table 2: Test performance in five benchmarks from Benchmarking GNNs (baselines please see Appx. A.5 for details and references). Shown is the mean $\pm$ s.d. of 4 runs with different random seeds. Highlighted are the top first, second, and third results. # Param under $500K$ for ZINC, PATTERN, CLUSTER and $\sim 100K$ for MNIST and CIFAR10.
@@
PGT (Ours) $\mathbf{0.0566\pm 0.002}$ $\mathbf{98.614\pm 0.096}$ $\mathbf{78.560\pm 0.700}$ $\mathbf{89.752\pm 0.030}$ $\mathbf{80.027\pm 0.114}$
@@
We present experimental results on two Long-range Graph Benchmark (LRGB) Dwivedi et al., 2022c datasets, Peptides-Function and Peptides-Structure in Table 5.2. PPGT achieves the lowest MAE on Peptides-Structure and remains in the top three models on Peptides-Function. Moreover, the difference in AP for Peptides-Function dataset between the best performing GRIT and our PPGT is not statistically significant at the 5% level for a one-tailed t-test. These results demonstrate PPGT’s outstanding capability of learning long-range dependency structures.
PCQM4Mv2 Large-scale Graph Regression Benchmark
Further, we conduct an experiment on the PCQM4Mv2 large-scale graph regression benchmark [^2] Clarification on large-scale graph-datasets and large-scale-graph datasets can be found in Appx. F.1. of 3.7M graphs Hu et al., (2021), which is currently one of the largest molecular datasets (Table 5.2). We follow the experimental setting of Rampášek et al., (2022) — excluding 3D information as attributes, and treating the validation set of PCQM4Mv2 as the test set, instead of the original private Test-dev set. The result of a single random seed run is reported due to the size of the dataset, following previous works. We followed the hyperparameter configuration of GRIT Ma et al., (2023) and did not conduct hyperparameter search due to computational limits. However, our model can still reach an outstanding performance, comparable to GRIT and GraphGPS, with similar parameter budget.
Table 3: Test performance on two benchmarks from Long-range Graph Benchmarks (LRGB). (mean $\pm$ s.d. of 4 runs). Highlighted are the top first, second, and third results.
Model Peptides-func Peptides-struct
[…]
GCN $0.5930\pm 0.0023$ $0.3496\pm 0.0013$
GINE $0.5498\pm 0.0079$ $0.3547\pm 0.0045$
GatedGCN+RWSE $0.6069\pm 0.0035$ $0.3357\pm 0.0006$
CKGCN ${0.6952\pm 0.0068}$ ${0.2477\pm 0.0018}$
Transformer+LapPE $0.6326\pm 0.0126$ $0.2529\pm 0.0016$
SAN+LapPE $0.6384\pm 0.0121$ $0.2683\pm 0.0043$
SAN+RWSE ${0.6439\pm 0.0075}$ $0.2545\pm 0.0012$
GPS $0.6535\pm 0.0041$ $0.2500\pm 0.0012$
GMLP-Mixer $\mathbf{0.6970\pm 0.0080}$ $\mathbf{0.2475\pm 0.0015}$
GRIT $\mathbf{0.6988\pm 0.0082}$ $\mathbf{0.2460\pm 0.0012}$
PPGT (Ours) $\mathbf{0.6961\pm 0.0062}$ $\mathbf{0.2450\pm 0.0017}$

### v3

@@ 3.4 Anti-Spectral-Bias with Sinusoidal PE Enhancement (SPE)
As discussed in Zhang et al. (2023b; 2024b), as the information provided through graph positional encodings (PEs) becomes more fine-grained, GD-GNNs can achieve better distinguishability of graph structure. However, due to the spectral bias of neural networks Rahaman et al. (2019), MLPs prioritize learning the low-frequency modes and thus lose detailed information stored in PEs.
To mitigate this issue, motivated by NeRF Mildenhall et al. (2020), we add an extra sinusoidal encoding on top of the RRWP and process it with a simple MLP. The sinusoidal encoding is applied to each channel of $\mathbf{p}^{\prime}_{ij}$ in an elementwise fashion:
@@
where $\mathtt{SPE}:\mathbb{R}\to\mathbb{R}^{1+2S}$ with $S\in\mathbb{Z}^{+}$ different bases. For notational conciseness, we use $\mathtt{SPE}(\mathbf{p}^{\prime}_{ij})\in\mathbb{R}^{K+2SK}$ to denote the application of $\mathtt{SPE}$ to all channels of $\mathbf{p}^{\prime}_{ij}$ and concatenate the outputs. The PE stem (Eq. (8)) becomes
@@ 3.5 Theoretical Expressivity of PPGT
The theoretical expressivity of PPGT can be analyzed within the GD-WL framework Zhang et al. (2023b). When equipped with regular generalized distances (e.g., RRWP, resistance distance Zhang et al. (2023b)) as graph PE, PPGT attains expressivity strictly beyond $1$-WL and is upper bounded by $3$-WL, matching the theoretical limits of a broad class of GD-WL algorithms. We provide the formal proof in Appx. E.1.
@@ 4 Relationship with Previous Work
Graph Transformers. Transformers, especially plain Transformers—architectures close to the vanilla Transformer Vaswani et al. (2017) with scaled dot-product (SDP) attention and feed-forward networks (FFNs)—have achieved outstanding performance across a wide range of domains, from language Devlin et al. (2019); Raffel et al. (2020); OpenAI (2024); Dubey et al. (2024) to vision Dosovitskiy et al. (2021); Touvron et al. (2021b); Caron et al. (2021); Oquab et al. (2024).
@@
Motivated by the success of Transformers in other domains, researchers have strived to migrate Transformers to graph learning to address the limitations of MPNNs. Although the naive migration of Transformers to graph learning did not work well Dwivedi and Bresson (2021), several recent works have achieved considerable success when constructing graph Transformers, from theoretical expressivity analysis Zhang et al. (2023b; 2024b) to impressive empirical performance Ying et al. (2021); Rampášek et al. (2022); Ma et al. (2023). However, during the development of these graph Transformers, there has been a gradual but substantial deviation from the plain Transformers widely used in other domains. For example, SAN Kreuzer et al. (2021) introduces dual-attention mechanisms with local and global aggregations; K-Subgraph SAT Chen et al. (2022) introduces MPNNs into attention mechanisms to compute attention scores; GraphGPS Rampášek et al. (2022) heavily relies on the MPNNs within its hybrid Transformer architecture; EGT Hussain et al. (2022) introduces gated mechanisms and edge-updates inside the attention; GRIT Ma et al. (2023) incorporates a complicated conditional MLP-based attention mechanism (shown in Appendix. D.1). These deviations prevent the easy adoption of Transformer training advances and obscure the potential unification of cross-modality foundation models. Among these graph Transformers, the graphormer-series Ying et al. (2021); Luo et al. (2022); Zhang et al. (2023b) retain architectures that are closest to plain Transformers; unfortunately, the result is a substantial gap in empirical performance and empirical expressivity compared to the best-performing graph Transformers Rampášek et al. (2022); Ma et al. (2023). Most graph Transformers treat a node as a token. TokenGT Kim et al. (2022) views both nodes and edges as tokens and processes them using plain or sparse Transformers. However, despite adopting a plain Transformer architecture, TokenGT also falls considerably behind recent graph Transformers on the expressivity-demanding PCQM4Mv2 dataset.
Developing powerful graph Transformers based on a plain Transformer architecture is particularly attractive, as the associated hardware stacks and software libraries have already been extensively optimized. Therefore, we investigate the fundamental limitations of plain Transformers on graph-structured data and introduce several simple yet effective architectural enhancements that preserve the core plain Transformer design, enabling competitive empirical performance without requiring significant architectural modifications.
@@
Other Attention Mechanisms. Besides SDP attention, there are other attention variants in use. Bahdanau et al. (2015) introduced the earliest content-based attention mechanism for recurrent neural networks (RNNs) based on an MLP, which is more computationally and memory costly. Swin-Transformer-V2 Liu et al. (2022) uses cosine-similarity to compute attention scores for better stability, but neglects the magnitude information. Kim et al. (2021) propose the use of the negative of the square $L_{2}$-distance, with tied query-key projection weights, for maintaining Lipchitz continuity of Transformers. Our attention mechanism, although based on $L_{2}$ attention, is further simplified and adjusted in order to maintain alignment with SDP attention.
Continuous Kernel Graph Convolution. Ma et al. (2024) introduce graph convolution operators with continuous kernels defined over pseudo-coordinates of graphs, termed CKGConv. These operators offer better flexibility in capturing high-frequency information compared to attention-based mechanisms. PPGT with URPE enhancement can also be interpreted as a generalization of CKGConv. In Eq. (5), the $\phi(\mathbf{p}_{ij})$ term serves as the convolution kernel, analogous to that in CKGConv, while the $\mathtt{Softmax}_{j}\big(\frac{\mathbf{q}_{i}^{\intercal}\mathbf{k}_{j}}{\sqrt{D}}-\frac{\mathbf{k}_{j}^{\intercal}\mathbf{k}_{j}}{2\sqrt{D}}+\theta(\mathbf{p}_{ij})\big)$ component can be regarded as a dynamic density function conditioned on the token representations, which is assumed to be uniform in CKGConv.
Universality of Transformers The Universal Approximation Theorem is fundamental for understanding the theoretical representational capacity upper bound of neural networks. While Hornik et al. (1989) established that MLPs are universal approximators for functions $f:\mathbb{R}^{N}\to\mathbb{R}^{M}$ on compact sets, this conclusion does not automatically extend to other domains. Yun et al. (2020) demonstrated that Transformers with absolute Positional Encodings (PE)—which sufficiently distinguish token order—are universal approximators for sequence-to-sequence functions. Conversely, Luo et al. (2022) showed that Transformers with relative PE lack this universality. However, they show that it can be restored through the universality enhancement. Regarding graph domains, Kreuzer et al. (2021) state that Transformers with absolute PE can approximate any function $f$ for the graph isomorphism problem. Crucially, however, this requires that the absolute PE can uniquely identify nodes in each graph, which is infeasible due to the highly symmetric structure of graphs.
Graph Transformers with full attention and appropriate positional encoding (PE) are typically considered at most as expressive as the 3-WL test for graph isomorphism Zhang et al. (2023b). In contrast, those using linear attention have expressivity equivalent to MPNNs with virtual nodes Cai et al. (2023).
@@
Additional related work. We discuss additional related work, including MPNNs, graph positional/structural encoding, higher-order GNNs, and subgraph GNNs in Appx. D.
@@ 5.1 Empirical Expressivity on Graph Isomorphism
To better understand the expressivity of PPGT, we evaluate our model on the BREC benchmark Wang and Zhang (2024), a comprehensive dataset for measuring the empirical expressive power of GNNs w.r.t. graph isomorphism, with graph-pairs from 1-WL to 4-WL-indistinguishable.
@@
[3]. Mismatch between expressivity and real-world benchmark performance: The stronger theoretical/empirical expressivity is not completely reflected in real-world benchmark performance. For example, subgraph GNNs and/or K-WL GNNs with stronger expressivity (e.g., SSWL+, I2GNN, N2GNN) demonstrate inferior performance compared to GRIT and PPGT on the ZINC benchmark.
[4]. Going beyond GD-WL?: As previously discussed, the design of graph PE, a.k.a., pseudo-coordinates, can be extended beyond the distance/affinity of graphs. With this in mind, we conduct an exploratory demo called I2GNN+PPGT, which uses I2GNN to generate additional positional encodings for PPGT. The empirical expressivity is further improved to 76%, outperforming the standalone I2GNN and PPGT, surpassing 3-WL and reaching the top performance among the methods compared. This demonstration hints that plain Graph Transformers can potentially surpass GD-WL and achieve greater expressive power purely through enhanced positional encoding designs. This result is noteworthy, as PPGT achieves superior empirical performance compared to many subgraph GNNs and higher-order GNNs that possess greater theoretical expressivity.
The expressivity bottleneck of our PPGT model does not stem from its architecture, but rather from the design of positional encodings (PE). To fully realize the expressive potential of plain Graph Transformers, it is crucial to develop expressive and generalizable PE schemes for graphs that are also compatible with permutation symmetry.
@@ 5.2 Benchmarking PPGT on Real-world Benchmarks
The specifications, details, and references for the baseline methods are provided in Appx. A.6.
@@
We conduct a general evaluation of our proposed PPGT on five datasets from Benchmarking GNNs (Dwivedi et al., 2022a): ZINC, MNIST, CIFAR10, PATTERN, and CLUSTER, and summarize the results in Table 2. We observe that our model obtains the best mean performance for all five datasets, outperforming various MPNNs, non-MPNN GNNs, and existing graph Transformers. These results showcase the effectiveness of PPGT for general graph learning with a plain Transformer architecture.
Table 2: Test performance on five benchmarks from Benchmarking GNNs (baselines please see Appx. A.6 for details and references). Shown is the mean $\pm$ s.d. of 4 runs with different random seeds. Highlighted are the top first, second, and third results. # Param under $500K$ for ZINC, PATTERN, CLUSTER and $\sim 100K$ for MNIST and CIFAR10.
@@
PPGT $\mathbf{0.0566\pm 0.002}$ $\mathbf{98.614\pm 0.096}$ $\mathbf{78.560\pm 0.700}$ $\mathbf{89.752\pm 0.030}$ $\mathbf{80.027\pm 0.114}$
@@
We present experimental results on three Long-range Graph Benchmark (LRGB) Dwivedi et al. (2022c) datasets – Peptides-Function, Peptides-Structure and PASCALVOC-SP in Table 3. PPGT achieves the lowest MAE on Peptides-Structure, top F1 on PascalVoc-SP and remains in the top three models on Peptides-Function. We adopt the updated experimental setup from LRGB as described in Tönshoff et al. (2024), and report the corresponding results. Moreover, the difference in AP for Peptides-Function dataset between the best performing GRIT and our PPGT is not statistically significant at the 5% level for a one-tailed t-test. These results demonstrate PPGT’s capability of learning long-range dependency structures.
5.3 Benchmarking PPGT on Large-scale Graph Benchmark and Large-scale-graph Benchmark
Scaling behavior with respect to both model size and dataset size is a critical consideration. Accordingly, in this section, we evaluate the applicability of PPGT to large-scale data settings without imposing a strict parameter budget.
We consider two distinct types of large-scale data settings of graphs: large-scale graph benchmarks and large-scale-graph benchmarks. Although prior work often conflates these two scenarios, the distinction is important.
Specifically, large-scale (graph) datasets comprise many graph instances, requiring scalability with respect to dataset size. In contrast, large-scale graph datasets consist of a single graph with an extremely large number of nodes, necessitating scalability with respect to input size (analogous to Gigapixel image processing). A more detailed discussion is provided in Appendix F.1.
Large-scale Graph Benchmark: PCQM4Mv2
To further assess the scalability of PPGT to large-scale data, we evaluate its performance on the PCQM4Mv2 large-scale graph regression benchmark, which consists of 3.7 million graphs Hu et al. (2021). This dataset is among the largest-scale graph benchmarks to date (see Table 5.3). Following the experimental protocol of Rampášek et al. (2022), we exclude the 3D information from the model attributes and use the PCQM4Mv2 validation set in place of the original private Test-dev set for evaluation. By omitting 3D information, we aim to more accurately measure each model’s intrinsic graph learning capability, ensuring that results do not reflect reliance on 3D coordinates.
The result of a single random seed run is reported due to the size of the dataset, following previous works. Our model performs well, comparably to GRIT and GraphGPS, with a similar parameter budget.
Table 3: Performance comparison on Long-Range Graph Benchmark (LRGB)Peptides and PascalVoc-SP datasets. (mean $\pm$ s.d. of 4 runs). Highlighted are the top first, second, and third results.
Method Peptides-Func Peptides-Struct PascalVoc-SP
AP $\uparrow$ MAE $\downarrow$ F1 $\uparrow$
GCN $0.6860\pm 0.0050$ $\mathbf{0.2460\pm 0.0007}$ $0.2078\pm 0.0031$
GINE $0.6621\pm 0.0067$ $\mathbf{0.2473\pm 0.0017}$ $0.2718\pm 0.0054$
GatedGCN $0.6765\pm 0.0047$ $0.2477\pm 0.0009$ $0.3880\pm 0.0040$
DRew $\mathbf{0.7150\pm 0.0044}$ $0.2536\pm 0.0015$ $0.3314\pm 0.0024$
Exphormer $0.6527\pm 0.0043$ $0.2481\pm 0.0007$ $\mathbf{0.3960\pm 0.0027}$
GPS $0.6534\pm 0.0090$ $0.2509\pm 0.0010$ $\mathbf{0.4440\pm 0.0065}$
GRIT $\mathbf{0.6988\pm 0.0082}$ $\mathbf{0.2460\pm 0.0012}$ -
PPGT $\mathbf{0.6961\pm 0.0062}$ $\mathbf{0.2450\pm 0.0017}$ $\mathbf{0.4641\pm 0.0033}$
@@ 5.3 Benchmarking PPGT on Large-scale Graph Benchmark and Large-scale-graph Benchmark
Large-scale-graph Benchmark: OGBN-ArXiv
Unlike large-scale graph benchmarks, which prioritize the model’s capacity to learn from many graph instances, a large-scale-graph benchmark typically emphasizes memory efficiency to handle the substantial size of an individual graph.
Even though PPGT is inherently not designed for processing large-size inputs, we also benchmark our approach on a large-scale-graph dataset: OGBN-ArXiv Hu et al. (2020) [^2] No feature enhancement techniques, such as UniMP Shi et al. (2021) or GIANT Chien et al. (2022), are applied..
To handle large graphs, we adopt an additional graph sampling strategy. Specifically, we convert node-level tasks into graph-level tasks by extracting a local inducted subgraph around each target node using breadth-first search (BFS) node sampling, referred to as Node2Subgraph conversion. Similar techniques exist Zeng et al. (2020); Sun et al. (2023), but we do not explore them here, as they are beyond the focus of this work. Even in efficiency-oriented graph Transformers, graph sampling techniques are widely adopted to handle ultra–large-scale input graphs that cannot be processed directly.
We compare PPGT with several state-of-the-art efficiency-oriented graph Transformers, including NodeFormer Wu et al. (2022), Exphormer Shirzad et al. (2023), and SGFormer Wu et al. (2023). These models aim for full node coverage on large graphs but often compromise expressivity, as linear-attention graph Transformers exhibit the equivalent theoretical expressivity as MPNNs with virtual nodes Cai et al. (2023). PPGT maintains high expressivity through Node2Subgraph sampling, even at the cost of reduced node coverage, and achieves performance comparable to these SOTA efficiency-oriented graph Transformers.
These results suggest that both model expressivity and node coverage are essential for learning on large graphs, and achieving a balance between the two is more important than overemphasizing either.
Table 5: Testing results (Accuracy) on OGBN-ArXiv (mean $\pm$ s.d. of 4 runs). Highlighted are the top first, second, and third results. Baseline results from  (Wu et al., 2023; Shirzad et al., 2023; Li et al., 2022). (*DeeperGCN does not provide s.d.)
Method GCN
GCN-
NSampler
DeeperGCN SIGN NodeFormer SGFormer Exphormer PPGT
OGBN-ArXiv $71.74\pm 0.29$ $68.50\pm 0.23$ $71.90$ $70.28\pm 0.25$ $59.90\pm 0.42$ $72.63\pm 0.13$ $72.44\pm 0.28$ $72.46\pm 0.15$